OpenStudy (anonymous):

solve system of equation 2x-2y-z=-6 x-5y+2z=-7 -5x+y+z=-21

jhonyy9 (jhonyy9):

2x-2y-z=-6 --- z=2x-2y+6 -5x+y+2x-2y+6=-21 --- -3x-y=-27 --- y=-3x+27 x-5y+2(2x-2y+6)=-7 --- x-5y+4x-4y+12=-7 --- 5x-9y=-19 --- 5x-9(-3x+27)=-19 --- 5x+27x-243=-19 --- 32x=224 --- x=224/32=7 if x=7 --- y=-3x+27=-21+27=6 so y=6 and z=2x-2y+6=14-12+6=8 x=7 y=6 z=8